You are a financial manager for Lenzie Corporation. The firm needs to purchase equipment for its new production plant. The total cost of the equipment is $1,500,000. It is estimated that the after-tax cash inflows from the project will be $710,000 annually for five years. Lenzie Corporation has a market value debt-to-assets ratio of 40%. The firm’s common stock has a beta of 1.85, the risk-free rate is 1.2%, and the expected return on the market portfolio is 7%. The firm's pre-tax cost of debt is 7%, and the flotation costs of equity and debt are 7% and 4%, respectively. The tax rate is 36%. Assume the project is of similar risk to the firm's existing operations. Based on the given information, please answer questions a, b, and c.

a) What is Lenzie Corporation's weighted average cost of capital?

b) What is Lenzie Corporation's weighted average floatation cost?

c) What is the NPV for the project after adjusting for flotation costs?
Answer #1

a)

Cost of equity = r(f) + Beta*(R(m) - r(f))

Cost of equity = 1.2% + 1.85*(7%-1.2%)

Cost of equity = 11.93%

Cost of debt = Pre-tax cost of debt*(1-tax rate)

Cost of debt = 7%*(1-36%) = 4.48%

WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.48%*0.4+11.93%*(1-0.4)
WACC =8.95%

b)

Weighted average floatation cost = Cost of debt*Floatation cost of debt + Cost of equity*Floatation cost of equity

Weighted average floatation cost = (1-0.4)*7% + 0.4*4%

Weighted average floatation cost = 5.8%

c)

NPV is found by discounting the cash-flows at WACC

Floatation costs are considered in the initial cash-flows

А B C 1 2 Year 3 0 Cash-flows =-1500000-5.8%*1500000 710000 710000 4 1 2 5 6 3 710000 7 4 710000 8 5 710000 9 10 NPV =C3+NPV(

A B с D 1 Year 2 3 0 4 1 2 5 6 Cash-flows -1587000 710000 710000 710000 710000 710000 3 7 4 8 5 9 10 NPV 1178238.918 11 12

Hence, NPV = $1178238.92